OpenStudy (anonymous):

what is the equation of the line in standard form that passes throught hte point (3 4) and is parallel to the line y=-x+6 a) x+y=1 b) x+y=7 c) x-y=-1 d) x-y=-7

OpenStudy (anonymous):

The answer is b. The gradient of the line y=-x+6 is -1, thus the gradient of line through (3 ,4) is also -1. Equation of a line is: y-y1=m(x-x1) substitute y1=4 m=-1 x1=3 so y-4=-1(x-3) rearrange to give y+x=7